OpenStudy (ria23):

Express the following function, F(x) as a composition of two functions f and g. F(x) = x^2/(x^2+6)

OpenStudy (anonymous):

you kind of have a lot of choices here

OpenStudy (anonymous):

you could day \[f(x)=\frac{x}{x+6}\] and \(g(x)=x^2\) making \[f(g(x))=f(x^2)=\frac{x^2}{x^2+6}\]

OpenStudy (anonymous):

but you have other options if you don't like that one

OpenStudy (ria23):

That works, I understand that one. ^.^ Thank yhu.